# Persistence of AR($1$) sequences with Rademacher innovations and linear mod $1$ transforms

### v2

@@ 6. Convergence of the conditional distributions
Thus, recalling that $Q_{a}=\{z:\varkappa_{a}(z)=\infty\}$, in either case we showed that if $z\not\in Q_{a}$, then $\nu_{a}((z-\varepsilon,z+\varepsilon))=0$ and the interval $(z-\varepsilon,z+\varepsilon)$ does not intersect $Q_{a}$. Hence the set $Q_{a}$ is closed and $\nu_{a}$ is supported on $Q_{a}$. As we have already mentioned in Section 5, for every $a\in(\frac{1}{2},\frac{2}{3})$, the invariant measure $\hat{\mu}_{a}$ of the transformation $\hat{T}_{a}$ is ergodic and is equivalent to the Lebesgue measure on $[0,1]$ by Theorem 3 in [17]. Hence for every $a\in(\frac{1}{2},\frac{2}{3})$, almost all orbits $\{\hat{T}_{a}^{k}(z):k\geq 0\}$ are normal. This implies that the set $Q_{a}$ has Lebesgue measure zero, and thus the measure $\nu_{a}$ is singular.
@@
as in the case of finite $\varkappa_{a}(z)$. We arrive at the same conclusion if the orbit of $z$ hits $1$ but does not hit $\frac{2a-1}{1-a}$, once we separately consider the points $z^{\prime}<z$ and $z^{\prime}\geq z$, as in the finite case; here $\varkappa_{a}(z^{\prime})$ and $\overline{F}_{a}(z^{\prime})$ are constant on $(z-\varepsilon,z)$ for some $\varepsilon>0$ and $\varkappa_{a}(z^{\prime})\to\infty$ as $z^{\prime}\to z+$. Likewise, we arrive at the same conclusion if the orbit of $z$ hits $\frac{2a-1}{1-a}$ but does not hit $1$; here $\varkappa_{a}(z^{\prime})$ and $\overline{F}_{a}(z^{\prime})$ are constant on $(z,z+\varepsilon)$ for some $\varepsilon>0$ and $\varkappa_{a}(z^{\prime})\to\infty$ as $z^{\prime}\to z-$ (if $z>0$). In each of the three cases, we have
@@
$\leq 2\sum_{j=N+1}^{\infty}\Big{(}\frac{p}{\lambda_{a}}\Big{)}^{j+1}\Big{(}\frac{q}{p}\Big{)}^{C(j+1)+1}$ (65)
with the constant $C>0$ as in Proposition 8 if $p<1/2$ and $C=0$ if $p\geq 1/2$. Taking $N\to\infty$ establishes continuity of $\overline{F}_{a}$ at point $z$.
It remains to consider the case where the orbit of $z$ hits both points $\frac{2a-1}{1-a}$ and $1$. This can only happen if $T_{a}^{k_{1}}(z)=1$ and $T_{a}^{k_{2}}(z)=\frac{2a-1}{1-a}$ for some $0\leq k_{1}<k_{2}$ (and hence $T_{a}^{k_{2}-k_{1}}(0)=\frac{1}{1-a})$. It is easy to see that in this case there is an $\varepsilon>0$ such that $\varkappa_{a}(z^{\prime})=k_{1}$ on $(z-\varepsilon,z)$; $\varkappa_{a}(z^{\prime})=k_{2}$ on $(z,z+\varepsilon)$; and $\overline{F}_{a}$ is constant on the whole of $(z-\varepsilon,z+\varepsilon)$ by (64). In this case $z$, which is an isolated point of $Q_{a}$, is not in the support of $\nu_{a}$. This completes the proof of continuity of $\overline{F}_{a}$ on the whole of the interval $[0,\frac{1}{1-a}]$.
It remains to show that the topological support of $\nu_{a}$ is the set $Q_{a}\setminus H_{a}$, where $H_{a}=\varnothing$ if $T^{k}_{a}(0)\neq\frac{1}{1-a}$ for all integer $1\leq k\leq\varkappa_{a}$, and $H_{a}=\cup_{k=0}^{\infty}T_{a}^{-k}(0)$ otherwise. So far we showed that $\nu_{a}$ is supported on $Q_{a}\setminus H_{a}$.
@@
Lastly, it is clear that $0=\overline{F}_{a}(\frac{1}{1-a})<\overline{F}_{a}(z^{\prime})$ for all $z^{\prime}<\frac{1}{1-a}$. Thus, we showed that if $a\in(\frac{1}{2},\frac{2}{3})$, then $\overline{F}_{a}$ is constant on no open neighbourhood of any point in the set $Q_{a}\setminus H_{a}$, which therefore is the topological support of $\nu_{a}$.
@@
The invariant measure $\hat{\mu}_{2/3}$ of the transformation $\hat{T}_{2/3}$ is ergodic and is equivalent to the Lebesgue measure on $[0,1]$ by Corollary to Theorem 2 in [17]. Then the measure $\mu_{2/3}$, defined by $\mu_{2/3}(A)=\hat{\mu}_{2/3}(A/3)$ for every Lebesgue measurable set $A\subset[0,3]$, is invariant and ergodic for $T_{2/3}$. Therefore, almost all orbits $\{T_{2/3}^{k}(z):k\geq 0\}$ are normal, and there exist a $z\in J$ and an $\varepsilon\in(0,1)$ such that the orbit of $z$ hits the interval $(1-\varepsilon,1)$ and $(z,z+\varepsilon)\subset J$. Denote by $k_{0}\geq 0$ the first hitting time. Since the mapping $T_{a}$ is piecewise continuous and satisfies $T_{a}^{\prime}=1/a>1$ on the interior of its domain, it follows that there exists a $z^{\prime}\in(z,z+\varepsilon)$ such that $T_{2/3}^{k_{0}}(z^{\prime})=1$. Then $\delta_{j}(z)=\delta_{j}(z^{\prime})$ for all $10\leq j\leq k_{0}-$ 1. Since $T_{2/3}(1-)=3$, it is easy to show, using representation (64) and arguing as above, that $\overline{F}_{2/3}(z)>\overline{F}_{2/3}(z^{\prime})$. Therefore, $\overline{F}_{2/3}$ is not constant on $J$, and thus $\nu_{2/3}(J)>0$, as needed.
3. Singularity properties of $\nu_{2/3}$.
In this part of the proof we assume throughout that $a=2/3$. To start, note that for the $\frac{3}{2}$-transformation $\bar{T}(x)=\frac{3}{2}x\pmod{1}$, we have
$T_{2/3}(x)=3(1-\bar{T}(1-x/3)),\qquad x\in[0,1)\cup(1,3].$
@@
Combined with (6), this result implies that for every $\delta\in(0,1)$ and almost every $x\in(0,3)$, there exist constants $\varepsilon_{\delta}(x)>0$ and $C_{\delta}(x)>0$ such that if $0<|\varepsilon|<\varepsilon_{\delta}(x)$, then
$|\overline{F}_{2/3}(x)-\overline{F}_{2/3}(x+\varepsilon)|\leq C_{\delta}(x)\left(\frac{p}{\lambda_{2/3}}\cdot\Big{(}\frac{q}{p}\Big{)}^{C}\right)^{(1-\delta)\log_{2/3}|\varepsilon|},$
where the constant $C>0$ is as in Proposition 8 if $p<1/2$ and $C=0$ if $p>1/2$. If $p\neq 1/2$, the base of the exponent on the r.h.s. above belongs to $(2/3,1)$ by (22). Therefore, by choosing $\delta$ to be small enough, we see that $|\overline{F}_{2/3}(x)-\overline{F}_{2/3}(x+\varepsilon)|=o(\varepsilon)$ for almost every $x$. Thus, $\overline{F}_{2/3}^{\prime}(x)=0$ for such $x$, which implies that $1-\overline{F}_{2/3}$ is a singular distribution function. It remains to prove the lemma.
@@
Let us obtain a matching lower bound. Notice that the density of the invariant measure $\mu_{2/3}$ of $T_{2/3}$, which is $\hat{h}_{2/3}(x/3)/3$, is bounded by (58). Combined with the Borel–Cantelli lemma, this implies that for $\mu_{2/3}$-almost every $x$, there exists an $n_{0}(x)\geq 1$ such that
@@
Let us take $k=\log_{2/3}|\varepsilon|+3\log_{2/3}\log_{2/3}|\varepsilon|$. For $30<|\varepsilon|<2/$ 3, we have $k>\log_{2/3}|\varepsilon|>1$ and
@@
4. The rate of convergence on test functions in BV.
Recall that $\operatorname*{Var}$ denotes the total variation of a function. We clam the following.
@@
Let $a\in(\frac{1}{2},\frac{2}{3}]$ and $p\in(0,1)$. Then there exist constants $C>0$ and $\gamma_{1}\in(0,1)$ such that for any function $f$ of bounded variation on $[0,1/(1-a)]$ and any $x\in[0,1/(1-a)]$,
[…]
@@
We extend the argument we gave above in Item 1 for $f_{z}$ to an arbitrary function $f$ of bounded variation. To this end, we estimate the spectral radius of the operator $P_{2}$. First,
@@
Therefore, $\|P_{2}^{n}\|\leq p^{n}$ if $p\geq 1/2$, otherwise we need a more delicate estimate. We present it without a proof (which takes another three pages).

### v3

@@ 6.2. Support and non-atomicity of $\nu_{a}$
Thus, recalling that $Q_{a}=\{z:\varkappa_{a}(z)=\infty\}$, in either case we showed that if $z\not\in Q_{a}$, then $\nu_{a}((z-\varepsilon,z+\varepsilon))=0$ and the interval $(z-\varepsilon,z+\varepsilon)$ does not intersect $Q_{a}$. Hence the set $Q_{a}$ is closed and $\nu_{a}$ is supported on $Q_{a}$. As we have already mentioned in Section 5, for every $a\in(\frac{1}{2},\frac{2}{3})$, the invariant measure $\widehat{\mu}_{a}$ of the transformation $\widehat{T}_{a}$ is ergodic and is equivalent to the Lebesgue measure on $[0,1]$ by Theorem 3 in [18]. Hence for every $a\in(\frac{1}{2},\frac{2}{3})$, almost all orbits $\{\widehat{T}_{a}^{k}(z):k\geq 0\}$ are normal. This implies that the set $Q_{a}$ has Lebesgue measure zero, and thus the measure $\nu_{a}$ is singular.
@@
as in the case of finite $\varkappa_{a}(z)$. We arrive at the same conclusion if the orbit of $z$ hits $1$ but does not hit $\frac{2a-1}{1-a}$, once we separately consider the points $z^{\prime}<z$ and $z^{\prime}\geq z$, as in the finite case; here $\varkappa_{a}(z^{\prime})$ and $\overline{F}_{a}(z^{\prime})$ are constant on $(z-\varepsilon,z)$ for some $\varepsilon>0$ and $\varkappa_{a}(z^{\prime})\to\infty$ as $z^{\prime}\to z+$. Likewise, we arrive at the same conclusion if the orbit of $z$ hits $\frac{2a-1}{1-a}$ but does not hit $1$; here $\varkappa_{a}(z^{\prime})$ and $\overline{F}_{a}(z^{\prime})$ are constant on $(z,z+\varepsilon)$ for some $\varepsilon>0$ and $\varkappa_{a}(z^{\prime})\to\infty$ as $z^{\prime}\to z$- if $z>0$ (the case $z=0$ is possible). In each of the three cases, we have
@@
$\leq 2r^{C}\sum_{j=N+1}^{\infty}\bigg(\frac{pr^{C}}{\lambda_{a}}\bigg)^{j+1}$ (77)
with $r=\max(1,q/p)$ and the constant $C>0$ as in Proposition 10. Taking $N\to\infty$ establishes continuity of $\overline{F}_{a}$ at point $z$.
It remains to consider the case where the orbit of $z$ hits both points $\frac{2a-1}{1-a}$ and $1$. This can only happen if $T_{a}^{k_{1}}(z)=1$ and $T_{a}^{k_{2}}(z)=\frac{2a-1}{1-a}$ for some $0\leq k_{1}<k_{2}$ (and hence $T_{a}^{k_{2}-k_{1}}(0)=\frac{1}{1-a})$. It is easy to see that in this case there is an $\varepsilon>0$ such that $\varkappa_{a}(z^{\prime})=k_{1}$ on $(z-\varepsilon,z)$; $\varkappa_{a}(z^{\prime})=k_{2}$ on $(z,z+\varepsilon)$; and $\overline{F}_{a}$ is constant on the whole of $(z-\varepsilon,z+\varepsilon)$ by (76). In this case $z$, which is an isolated point of $Q_{a}$, is not in the support of $\nu_{a}$. This completes the proof of continuity of $\overline{F}_{a}$ on the whole of the interval $0,\frac{1}{1-a}]$.
It remains to show that the topological support of $\nu_{a}$ is the set $Q_{a}\setminus H_{a}$, where $H_{a}=\varnothing$ if $T^{k}_{a}(0)\neq\frac{1}{1-a}$ for all integer $1\leq k\leq\varkappa_{a}$, and $H_{a}=\cup_{k=0}^{\infty}T_{a}^{-k}(0)$ otherwise. Our proof above of the continuity of $\overline{F}_{a}$ at points $z$ with $\varkappa_{a}(z)=\infty$ actually showed that $\nu_{a}$ is supported on $Q_{a}\setminus H_{a}$, and that each $z\in H_{a}$ is an isolated point of $Q_{a}$.
@@
Lastly, it is clear that $0=\overline{F}_{a}(\frac{1}{1-a})<\overline{F}_{a}(z^{\prime})$ for all $z^{\prime}<\frac{1}{1-a}$. Thus, we showed that if $a\in(\frac{1}{2},\frac{2}{3})$, then $\overline{F}_{a}$ is constant on no open neighbourhood of any point in the set $Q_{a}\setminus H_{a}$, which therefore is the topological support of $\nu_{a}$. It has no isolated points since $\overline{F}_{a}$ is continuous.
@@
The invariant measure $\widehat{\mu}_{2/3}$ of the transformation $\widehat{T}_{2/3}$ is ergodic and is equivalent to the Lebesgue measure on $[0,1]$ by Corollary to Theorem 2 in [18]. Then the measure $\mu_{2/3}$, defined by $\mu_{2/3}(A)=\widehat{\mu}_{2/3}(A/3)$ for every Lebesgue measurable set $A\subset[0,3]$, is invariant and ergodic for $T_{2/3}$. Therefore, almost all orbits $\{T_{2/3}^{k}(z):k\geq 0\}$ are normal, and there exist a $z\in J$ and an $\varepsilon\in(0,1)$ such that the orbit of $z$ hits the interval $(1-\varepsilon,1)$ and $(z,z+\varepsilon)\subset J$. Denote by $k_{0}\geq 0$ the first hitting time. Since the mapping $T_{a}$ is piecewise continuous and satisfies $T_{a}^{\prime}=1/a>1$ on the interior of its domain, it follows that there exists a $z^{\prime}\in(z,z+\varepsilon)$ such that $T_{2/3}^{k_{0}}(z^{\prime})=1$. Then $\delta_{j}(z)=\delta_{j}(z^{\prime})$ for all $0\leq j\leq k_{0}-1$. Since $T_{2/3}(1-)=3$, it is easy to show, using representation (76) and arguing as above, that $\overline{F}_{2/3}(z)>\overline{F}_{2/3}(z^{\prime})$. Therefore, $\overline{F}_{2/3}$ is not constant on $J$, and thus $\nu_{2/3}(J)>0$, as needed.
6.3. Singularity properties of $\nu_{2/3}$
In this part of the proof we assume throughout that $a=2/3$. To start, note that for the $\frac{3}{2}$-transformation $\overline{T}(x)=\frac{3}{2}x\pmod{1}$, we have
$T_{2/3}(x)=3(1-\overline{T}(1-x/3)),\qquad x\in[0,1)\cup(1,3].$
@@ 6.3. Singularity properties of $\nu_{2/3}$
Combined with (6.2), this result implies that for every $\delta\in(0,1)$ and almost every $x\in(0,3)$, there exist constants $\varepsilon_{\delta}(x)>0$ and $C_{\delta}(x)>0$ such that if $0<|\varepsilon|<\varepsilon_{\delta}(x)$, then
$|\overline{F}_{2/3}(x)-\overline{F}_{2/3}(x+\varepsilon)|\leq C_{\delta}(x)\left(\frac{pr^{C}}{\lambda_{2/3}}\right)^{(1-\delta)\log_{2/3}|\varepsilon|},$
where $r=\max(1,q/r)$ and the constant $C>0$ is as in Proposition 10. If $p\neq 1/2$, then $pr^{C}/\lambda_{2/3}\in(2/3,1)$ by (31). Therefore, by choosing $\delta$ to be small enough, we see that $|\overline{F}_{2/3}(x)-\overline{F}_{2/3}(x+\varepsilon)|=o(\varepsilon)$ for almost every $x$. Thus, $\overline{F}_{2/3}^{\prime}(x)=0$ for such $x$, which implies that $1-\overline{F}_{2/3}$ is a singular distribution function. It remains to prove the lemma.
@@
Let us obtain a matching lower bound. Notice that the density of the invariant measure $\mu_{2/3}$ of $T_{2/3}$, which is $\widehat{h}_{2/3}(x/3)/3$, is bounded by (68). Combined with the Borel–Cantelli lemma, this implies that for $\mu_{2/3}$-almost every $x$, there exists an $n_{0}(x)\geq 1$ such that
@@
Let us take $k=\log_{2/3}|\varepsilon|+3\log_{2/3}\log_{2/3}|\varepsilon|$. For $0<|\varepsilon|<2/3$, we have $k>\log_{2/3}|\varepsilon|>1$ and
@@
6.4. Rate of convergence on test functions in BV
Recall that $\operatorname*{Var}[f]$ denotes the total variation of a function $f$. We clam the following.
@@ 6.4. Rate of convergence on test functions in BV
Let $a\in(\frac{1}{2},\frac{2}{3}]$ and $p\in(0,1)$. Then there exist constants $C_{3}>0$ and $\gamma_{1}\in(0,1)$ such that for every function $f$ of bounded variation on [0,1/(1a)][0,1/(1-a)], $x\in[0,1/(1-a)]$, and $n\geq 1$, we have
$\bigg|\mathbb{E}_{x}(f(X_{n})|\tau>n)-\int_{[0,\frac{1}{1-a}]}fd\nu_{a}\bigg|\leq C_{3}\gamma_{1}^{n}\operatorname*{Var}[f].$ (82)
@@
We extend the argument we gave above in Section 6.1 for $f_{z}$ to an arbitrary function $f$ of bounded variation. To this end, we estimate the spectral radius of the operator $P_{2}$. First,
@@
Therefore, $\|P_{2}^{n}\|\leq p^{n}$ if $p\geq 1/2$, otherwise we need a more delicate estimate below.
@@
We postpone the proof of this estimate and first finish the proof of Proposition 18; note that the right-hand side of (83) equals $p^{n}$ when $p\geq 1/2$.
Remark 20.
Proposition 19 combined with the estimates of Proposition 10 imply that the spectral radius of $P_{2}$ on BVBV satisfies $\rho(P_{2})<\lambda_{a}$. This implies that the operator $P$ is quasi-compact on $BV$ for all $a\in(\frac{1}{2},\frac{2}{3}]$, since $\rho(P)\geq\lambda_{a}$ and the operator $P_{1}=P-P_{2}$ is compact because its range is two-dimensional; cf. Remark 15.